# Think2Drive: Efficient Reinforcement Learning by Thinking in Latent World Model for Quasi-Realistic Autonomous Driving (in CARLA-v2)

### v1

@@ 4.1 CARLA Leaderboard v2
CARLA Leaderboard v2 is based on the CARLA simulator with version bigger than 0.9.13 (V1 on 0.9.10). We evaluate our planner with CARLA 0.9.14. Along with their simulator, CARLA team initially proposed Leaderboard v1, which is composed of basic tasks such as lane following, turning, collision avoidance, and etc. Then, to facilitate quasi-realistic urban driving, they propose CARLA v2, which encompasses a multitude of complex scenarios previously absent in v1. These scenarios pose serious challenges to the ability of autonomous driving methods.
[…]
4.2 CornerCaseRepository Benchmark
As mentioned below, in the official benchmark, multiple scenarios are along a single long route, which makes it hard to train and evaluate the model. To address this deficiency, we introduce the CornerCaseRepository benchmark. It consists of 1600 routes for training and 390 routes for evaluation. Every route in CornerCaseRepository benchmark contains only one single scenario with a length $<$ 300 meters so that the training and evaluation of different scenarios are decoupled. In the training routes, there are 40 routes for each scenario and 40 routes without any scenarios. The routes are sampled randomly during the RL training process. There are 10 routes for each scenario in evaluation routes. In evaluation mode, the routes are sampled sequentially until all routes have been evaluated. The CornerCaseRepository benchmark supports the use of the CARLA metrics (e.g. driving scores, route completion) to analyze the performance of each scenario separately, which provides convenience for debugging.
@@ 4.3 Weighted Driving Score
As described in Sec. 4.1, the scoring rules of CARLA leaderboard is too rough for driving policy evaluation. For instance, consider a driving model with an average infraction rate of 0.2 per kilometer and a penalty factor of 0.8. Under the hypothetical ideal condition where route completion is 100% for both 5-kilometer and 10-kilometer test routes, the driving scores would be 0.8 and 0.64, i.e., the longer the distance traveled, the lower the final driving score. To avoid such counter-intuitive phenomenon, we propose a new metric named Weighted Driving Score (WDS), formed as:
@@
where RC means route completion rate, $m$ is the total number of types of infractions considered, $\text{penalty}_{i}$ refers to the penalty factor for infraction type $i$ officially defined by CARLA team, and $n_{i}=\frac{\text{Number of Infractions}}{\text{Number of Scenarios}}$ (when there is no scenario, we simply set $n_{i}=\text{Number of Infractions}$ and in this case Weighted Driving Score=Driving Score.). Weighted Driving Score effectively balances the weight between route completion, the number of infractions, and scenario density, which provides a perspective about the average infractions when agent encoutering scenarios.
@@ 4.4 Performance
Tab. 1 and Fig. 1 gives results on the CornerCaseRepository benchmark. The overall training time on one A6000 GPU with AMD Epyc 7542 CPU - 128 logical cores is 3 days. For the baseline expert model, we implement Roach [38], where we replace our model-based RL model with model-free PPO [31] and keep all the others the same. Both experts are trained on 1600 routes and evaluated on other 390 routes of CornerCaseRepository benchmark for 3 runs. Think2Drive outperforms Roach by a large margin, showing the advantages of model-based RL.
@@
Figure 1: Driving performance and infractions on CornerCaseRepository benchmark.
We also choose and train an end-to-end baseline namely TCP [35], a lightweight student model with a strong performance on CARLA Leaderboard v1, as the imitation learning agent. We train TCP with 200K frames collected by the Think2Drive expert under different weather. We could observe that TCP, as a student model with only raw sensor inputs, has a large performance gap with both expert models, as caused by the difficulty of perception parts as well as the imitation learning process.
We also evaluate Think2Drive on the official test routes of CARLA v1 & v2 and compare it with the expert model of Roach. As illustrated in Tab. 3, Think2Drive not only outperforms Roach in the easier CARLA v1(no scenarios) but also achieves significantly superior performance in the more complex v2 routes. PPO achieved remarkably low scores, a consequence of its rapid convergence to local optima, beyond which the policy ceases to improve further even with the help of the reset technique. We argue that the reason for this phenomenon is that after each reset, PPO has to relearn the policy from the trajectories stored in the replay buffer, which contain inherent reward noise due to the AD characteristics. Conversely, a model-based planner can get accurate and smooth rewards from the world model.
Figure 2: Example of prediction by the world model based on the first 5 frames. The model can predict reasonable future frames. In the failure case, the planner runs a red light and the world model terminates the episode, so the following predictions are randomly generated.
Table 3: Performance on official test routes.
Method benchmark
Driving
Scores
Weighted
Driving Score
Route
Complete %
Roach (Expert) CARLA v1 85.0 72.0
Think2Drive (Ours) 90.2 90.2 99.7
PPO (Expert) CARLA v2 0.7 0.6 1.0
Think2Drive (Ours) 56.8 91.7 98.6
@@ 4.5 Infraction Analysis on Hard Scenarios
The scenarios Accident, Construction, and HazardAtSidelane along with their respective TwoWays versions, belong to the category of RouteObstacles scenarios. In these scenarios, the ego vehicle is required to perform lane changes to maneuver around obstacles, particularly in the case of the TwoWays versions where the ego vehicle needs to switch to the opposite lane. Such scenarios demand the ego vehicle to acquire a sophisticated lane negotiation policy, especially in the TwoWays scenarios where the ego vehicle must execute lane changing, maneuver around obstacles, and return to its original lane within a short time window. Failed cases in these scenarios typically result from collisions with an opposite car during the process of returning to the original lane after bypassing the obstacles. CARLA v2 generates randomly the opposite traffic flow with speed and interval range within [8,18][8,18][ 8 , 18 ] and [15,50][15,50][ 15 , 50 ] (typical value, may vary with specific road conditions) in the TwoWays scenarios, which may lead to a significantly constrained time window (e.g.less 1 second) for bypassing obstacles when the speed is large while the interval is small. Consequently, the ego vehicle is required to rapidly accelerate from $speed=0$ to its maximum speed, and the ego vehicle usually runs at a high speed and is close to the opposite car when returning to its original lane, which leads to a high risk of collisions.
The scenarios SignalizedLeftTurn, CrossingBicycleFlow, SignalizedRightTurn and BlockedInterSection belong to JunctionNegotiate type. In these scenarios, the ego vehicle has to interrupt the opposite dense car or bicycle flow, merge into the dense traffic flow, and stop at the junction to await road clearance. In CARLA v2, the traffic flow of these scenarios is configured to be very aggressive, meaning it does not proactively yield to the ego vehicle. The ego vehicle needs to maintain a reasonable distance from other vehicles to avoid collisions. For instance, in the SignalizedRightTurn scenario, the ego vehicle is expected to merge into traffic with an interval within [15,25][15,25][ 15 , 25 ] meters and a speed within [12,20][12,20][ 12 , 20 ] m/s. With a vehicle length of approximately 3 meters, the ego vehicle must not only accelerate rapidly to match the traffic speed in a short time but also simultaneously maintain a safe following distance from preceding and succeeding vehicles.
In scenario EnterActorFlow, when the ego vehicle enters the junction, a flow of cars from another junction entry runs a red light, which either will go for the same junction exit to the ego or intersect with the ego’s target route. The speed of this car flow is configured at a high level, ranging between 15 and 20 meters per second. At such high speeds, even minor steering inputs can result in significant changes in the vehicle’s heading, making it challenging for the ego vehicle to maintain frontal stability which significantly increases the risk of collisions.
@@ 4.6 Visualization of World Model Prediction
Throt. Brake Steer Throt. Brake Steer Throt. Brake Steer
@@
4.7 Ablation Study
Figure 3: Ablation of different bricks devised in the paper.
5 Conclusion
We have proposed a purely learning-based planner, Think2Drive, for quasi-realistic traffic scenarios. Benefiting from the model-based RL paradigm, It can drive proficiently in CARLA v2 with all 39 scenarios within 3 days of training on a single GPU. We also devise tailored bricks such as resetting technique, automated scenario generation, termination-priority replay strategy, and steering cost function to address the obstacles associated with applying model-based RL to autonomous driving tasks. Think2Drive highlights and validates a feasible approach, model-based RL, for quasi-realistic autonomous driving. Our model can also serve as a data collection model, providing expert driving data for end-to-end autonomous driving models.
$C=34$
[…]
road route ego lane
yellow
line
white
line
vehicle walker
emergency
car
obstacle
green
traffic light
yellow&red
traffic light
stop
sign
Table 6: Dimension of input/output representation. The $T$ temporal masks are from historical time-steps: [16,11,6,1][-16,-11,-6,-1][ - 16 , - 11 , - 6 , - 1 ].
H W C T M
128 128 34 4 30
6 Implementation Details
6.1 Input&Output Representation.
For the input representation, we utilize BEV semantic segmentation masks $i_{RL}\in\{0,1\}^{H\times W\times C}$ as image input, where each channel denotes the occurrence of certain types of objects. It is generated from the privileged information obtained from the simulator and consists of $C$ masks of size $H\times W$. In these $C$ masks, the route, lanes, and lane markings are all static and thus could be represented by a single mask while those dynamic objects (e.g. vehicles and pedestrians) have $T$ masks where each mask represents their state at one history time-step. Additionally, we feed speed, control action, and relative height of the ego vehicle at the previous time steps as input $v_{RL}\in\mathbb{R}^{K}$.
For output representation, we discretize the continuous action spaces of $throttle\in[0,1]$, $brake\in[0,1]$, and $steer\in[0,1]$ into $M$ actions to reduce the complexity. The actor outputs a one-hot action vector $a\in\{0,1\}^{M}$. The discretized actions are presented in Tab. 4
The aforementioned hyper-parameters are set as in Tab. 6 while the details about the input information are in Tab. 5.
6.2 Simulator Execution Mode
We boost its running efficiency via asynchronous reloading and parallel execution which could avoid the waiting time caused by the long preparation time of starting a new route, as shown in Fig. 4. For each 100K steps of execution, it can reduce about 1 day time cost.

### v2

@@ 4.1 CARLA Leaderboard v2
CARLA v2 is based on the CARLA simulator with version bigger than 0.9.13 (V1 on 0.9.10). We evaluate our planner with CARLA 0.9.14. CARLA team initially proposed Leaderboard v1, which is composed of basic tasks such as lane following, turning, collision avoidance, and etc. Then, to facilitate quasi-realistic urban driving, CARLA v2 is released, which encompasses multitude complex scenarios previously absent in v1. These scenarios pose serious challenges.
Since the release of CARLA v2, no team has managed to get a spot to tackle these scenarios, despite the availability of perfect logs scoring 100% on each scenario, provided by the CARLA official platform to aid in related research. In our analysis, there are four primary reasons to the difficulty of v2: 1) Extended Route Lengths: In CARLA v2, the routes extend between 7 to 10 kilometers, a substantial increase from the roughly 1-kilometer routes in v1. 2) Complex and Abundant Scenarios: Each route contains around 60 scenarios, which require the driving methods to be able to handle complex road conditions and conduct subtle control. 3) Exponential decay scoring rules: The leaderboard employs a scoring mechanism that penalizes infractions through multiplication penalty factors $<1$. In scenarios with extended routes and a multitude of scenarios, models struggle to attain high scores. 4) Limited data: the CARLA team only provides a set of 90 training routes coupled with scenarios while routes randomly generated by researchers, does not have official API support for the placement of scenarios.
4.2 CornerCaseRepo Benchmark
In the official benchmark, multiple scenarios are along a single long route, making it hard to train and evaluate the model. To address this deficiency, we introduce the CornerCaseRepo benchmark, consisting of 1,600 routes for training and 390 routes for evaluation. Every route in the benchmark contains only one type of scenario with a length $<$ 300 meters so that the training and evaluation of different scenarios are decoupled. In the training set, there are 40 routes for each scenario and 40 routes without any scenarios. The routes are sampled randomly during the RL training process. There are 10 routes for each scenario in the evaluation routes. For evaluation, the routes are sampled sequentially until all routes have been evaluated. CornerCaseRepo supports the use of the CARLA metrics (e.g. driving scores, route completion) to analyze the performance of each scenario separately, providing convenience for debugging.
@@ 4.3 Weighted Driving Score
As described in Sec. 4.1, the scoring rules of CARLA leaderboard is imperfect for driving policy evaluation. For instance, consider a driving model with an average infraction rate of 0.2 per kilometer and a penalty factor of 0.8. Under the hypothetical ideal condition where route completion is 100% for both 5-kilometer and 10-kilometer test routes, the driving scores would be 0.8 and 0.64, i.e., the longer the distance traveled, the lower the final driving score. To avoid such counter-intuitive phenomenon, we propose a new metric named Weighted Driving Score (WDS), formed as:
@@
where RC means route completion rate, $m$ is the total number of types of infractions considered, $\text{penalty}_{i}$ is the penalty factor for infraction type $i$ officially defined in CARLA, and $n_{i}=\frac{\text{Number of Infractions}}{\text{Scenario Density}}$ (when there is no scenario, we set $n_{i}=\text{Number of Infractions}$ in which case Weighted Driving Score=Driving Score). Weighted Driving Score effectively balances the weight between route completion, number of infractions, and scenario density, providing a measure of the average infractions encountered by the ego vehicle over routes.
Table 3: Performance on official test routes.
Method benchmark Driving Scores Weighted Driving Score Route Complete %
Roach (Expert) CARLA Leaderboard v1 84.0 95.0
Think2Drive (Ours) 90.2 90.2 99.7
PPO (Expert) CARLA Leaderboard v2 0.7 0.6 1.0
Think2Drive (Ours) 56.8 91.7 98.6
@@ 4.4 Performance
Tab. 1 and Fig. 1(a) show results on CornerCaseRepoṪhe overall training time on one A6000 GPU with AMD Epyc 7542 CPU – 128 logical cores is 3 days. For the baseline expert model, we implement Roach [45], where we replace our model-based RL model with model-free PPO [36] and keep all other techniques the same. Both experts are trained on 1600 routes and evaluated on other 390 routes of CornerCaseRepo benchmark for 3 runs. Think2Drive outperforms Roach by a large margin, showing the advantages of model-based RL.
@@
Figure 1: Driving performance and infractions on CornerCaseRepo.
We also choose and train an end-to-end baseline TCP [42], a lightweight yet competitive student model on CARLA Leaderboard v1, as the imitation learning agent. We train TCP with 200K frames collected by the Think2Drive expert under different weather. We could observe that TCP, as a student model with only raw sensor inputs, has a large performance gap with both expert models, as caused by the difficulty of perception as well as the imitation process.
We also evaluate Think2Drive on the official test routes of CARLA Leaderboard v1 & v2 and compare it with the expert model of Roach. As illustrated in Tab. 3, Think2Drive not only outperforms Roach in the easier CARLA Leaderboard v1(no scenarios) but also achieves significantly superior performance in the more complex v2 routes. PPO achieved remarkably low scores, a consequence of its rapid convergence to local optima, beyond which the policy ceases to improve further even with the help of the reset technique. We argue that the reason for this phenomenon is that after each reset, PPO has to relearn the policy from the trajectories stored in the replay buffer, which contain inherent reward noise due to the AD characteristics. Conversely, a model-based planner can get accurate and smooth rewards from the world model.
@@ 4.5 Infraction Analysis on Hard Scenarios
The scenarios Accident, Construction, and HazardAtSidelane along with their respective TwoWays versions, belong to the category of RouteObstacles scenarios. In these scenarios, the ego vehicle is required to perform lane changes to maneuver around obstacles, particularly in the case of the TwoWays versions where the ego vehicle needs to switch to the opposite lane. Such scenarios demand the ego vehicle to acquire a sophisticated lane negotiation policy, especially in the TwoWays scenarios where the ego vehicle must execute lane changing, maneuver around obstacles, and return to its original lane within a short time window. Failed cases in these scenarios typically result from collisions with an opposite car during the process of returning to the original lane after bypassing the obstacles. CARLA Leaderboard v2 generates randomly the opposite traffic flow with speed and interval range within $[8,18]$ and $[15,50]$ (typical value, may vary with specific road conditions) in the TwoWays scenarios, which may lead to a significantly constrained time window (e.g.less 1 second) for bypassing obstacles when the speed is large while the interval is small. Consequently, the ego vehicle is required to rapidly accelerate from $speed=0$ to its maximum speed, and the ego vehicle usually runs at a high speed and is close to the opposite car when returning to its original lane, which leads to a high risk of collisions.
The scenarios SignalizedLeftTurn, CrossingBicycleFlow, SignalizedRightTurn and BlockedInterSection belong to JunctionNegotiate type. In these scenarios, the ego vehicle has to interrupt the opposite dense car or bicycle flow, merge into the dense traffic flow, and stop at the junction to await road clearance. In CARLA Leaderboard v2, the traffic flow of these scenarios is configured to be very aggressive, meaning it does not proactively yield to the ego vehicle. The ego vehicle needs to maintain a reasonable distance from other vehicles to avoid collisions. For instance, in the SignalizedRightTurn scenario, it is expected to merge into traffic with an interval within $[15,25]$ meters and a speed within $[12,20]$ m/s. With a vehicle length of approximately 3 meters, the ego vehicle must not only accelerate rapidly to match the traffic speed in a short time but also maintain a safe following distance from other vehicles.
Figure 2: Prediction by the world model using the first 5 frames. It can predict reasonable future frames. In the failure case, the planner runs a red light and the world model terminates the episode, so the subsequent predictions are randomly generated.
@@ 4.6 Visualization of World Model Prediction
Figure 3: Ablation of different bricks devised in the paper.
4.7 Ablation Study
5 Implement Details
For the input representation, we utilize BEV semantic segmentation masks $i_{RL}\in\{0,1\}^{H\times W\times C}$ as image input, where each channel denotes the occurrence of certain types of objects. It is generated from the privileged information obtained from the simulator and consists of $C$ masks of size $H\times W$. In these $C$ masks, the route, lanes, and lane markings are all static and thus could be represented by a single mask while those dynamic objects (e.g. vehicles and pedestrians) have $T$ masks where each mask represents their state at one history time-step. Additionally, we feed speed, control action, and relative height of the ego vehicle at the previous time steps as input $v_{RL}\in\mathbb{R}^{K}$. More specific details in Appendix 0.B, Tab. 4 and Tab. 5.
For output representation, we discretize the continuous action into 30 actions to reduce the complexity. The discretized actions are presented in Tab. 6.
Reward Shaping. It is shaped to make the planner keep safe driving and finish the route as much as possible, which consists of four parts: 1) Speed reward $r_{speed}$ is used to train the ego vehicle to keep a safe speed, depending on the distance to other objects and their type. 2) Travel reward $r_{travel}$ is the distance traveled along the target routes at each tick of CARLA. Travel reward encourages the ego vehicle to finish more target routes. 3) Deviation penalty $p_{deviation}$ is the negative value of the distance between the ego vehicle and the lane center. It is normalized by the max deviation threshold $D_{max}$. 4) Steering cost $c_{steer}$ is used to make the ego vehicle drive more smoother. We set it as the difference between the current steer and the last one. The overall reward is given by:
$r=r_{speed}+\alpha_{tr}r_{travel}+\alpha_{de}p_{deviation}+\alpha_{st}c_{steer}$ (8)
6 Conclusion
We have proposed a purely learning-based planner, Think2Drive, for quasi-realistic traffic scenarios. Benefiting from the model-based RL paradigm, It can drive proficiently in CARLA Leaderboard v2 with all 39 scenarios within 3 days of training on a single GPU. We also devise tailored bricks such as resetting technique, automated scenario generation, termination-priority replay strategy, and steering cost function to address the obstacles associated with applying model-based RL to autonomous driving tasks. Think2Drive highlights and validates a feasible approach, model-based RL, for quasi-realistic autonomous driving. Our model can also serve as a data collection model, providing expert driving data for end-to-end autonomous driving models.
Appendix 0.A Discription of CARLA v2 Scenarios
CARLA v2 provides 39 corner scenarios which are common in real driving environments. These scenarios can generally be divided into two categories: regular road scenarios and junction scenarios. Here we give a detailed description of each scenario:
0.A.1 Regular road scenarios
These scenarios are related to regular road segments
1. ControlLoss
The ego vehicle loses control due to bad conditions on the road and it must recover, coming back to its original lane.
2. ParkingExit
The ego vehicle must exit a parallel parking bay into a flow of traffic.
3. ParkingCutIn
The ego vehicle must slow down or brake to allow a parked vehicle exiting a parallel parking bay to cut in front.
4. StaticCutIn
The ego vehicle must slow down or brake to allow a vehicle of the slow traffic flow in the adjacent lane to cut in front. Compared to ParkingCutIn, there are more cars in the adjacent lane and any one of them may cut in.
5. ParkedObstacle
The ego vehicle encounters a parked vehicle blocking part of the lane and must perform a lane change into traffic moving in the same direction to avoid it.
6. ParkedObstacleTwoWays
The ’TwoWays’ version of ParkedObstacle. The ego vehicle encounters a parked vehicle blocking the lane and must perform a lane change into traffic moving in the opposite direction to avoid it.
7. Construction
The ego vehicle encounters a construction site blocking and must perform a lane change into traffic moving in the same direction to avoid it. Compared to ParkedObstacle, the construction occupies more width of the lane. The ego vehicle has to completely deviate from its task route temporarily to bypass the construction zone.
8. ConstructionTwoWays
The ’TwoWays’ version of Construction.
9. Accident
The ego vehicle encounters multiple accident cars blocking part of the lane and must perform a lane change into traffic moving in the same direction to avoid it. Compared to ParkedObstacle and Construction, these accident cars occupy more length along the lane. The ego vehicle has to completely deviate from its task route for a longer time to bypass the accident zone.
10. AccidentTwoWays
The ’TwoWays’ version of Accident. Compared to ParkedObstacleTwoWays and ConstructionTwoWays, there is a much shorter time window for the ego vehicle to bypass the route obstacles (i.g. accident cars).
11. HazardAtSideLane
The ego vehicle encounters a slow-moving hazard blocking part of the lane. The ego vehicle must brake or maneuver next to a lane of traffic moving in the same direction to avoid it.
12. HazardAtSideLaneTwoWays
The ego vehicle encounters a slow-moving hazard blocking part of the lane. The ego vehicle must brake or maneuver to avoid it next to a lane of traffic moving in the opposite direction.
13. VehiclesDooropenTwoWays
The ego vehicle encounters a parked vehicle opening a door into its lane and must maneuver to avoid it.
14. DynamicObjectCrossing
A walker or bicycle behind a static prop crosses the road suddenly when the ego vehicle is close to the prop. The ego vehicle must make a hard brake promptly.
15. ParkingCrossingPedestrian
The ego vehicle encounters a pedestrian emerging from behind a parked vehicle and advancing into the lane. The ego vehicle must brake or maneuver to avoid it. Compared to DynamicObjectCrossing, the pedestrian is closer to the road and the ego vehicle has to act more timely.
16. HardBrake
The leading vehicle decelerates suddenly and the ego vehicle must perform an emergency brake or an avoidance maneuver.
17. YieldToEmergencyVehicle
The ego vehicle is approached by an emergency vehicle coming from behind. The ego vehicle must maneuver to allow the emergency vehicle to pass.
18. InvadingTurn
When the ego vehicle is about to turn right, a vehicle coming from the opposite lane invades the ego’s lane, forcing the ego to move right to avoid a possible collision.
0.A.2 Junction scenarios
These scenarios are related to junctions.
1. PedestrainCrossing
While the ego vehicle is entering a junction, a group of natural pedestrians suddenly cross the road and ignore the traffic light. The ego vehicle must stop and wait for all pedestrians to pass even though there is a green traffic light or a clear junction.
2. VehicleTurningRoutePedestrian
While performing a maneuver, the ego vehicle encounters a pedestrian crossing the road and must perform an emergency brake or an avoidance maneuver.
3. VehicleTurningRoute While performing a maneuver, the ego vehicle encounters a bicycle crossing the road and must perform an emergency brake or an avoidance maneuver. Compared to VehicleTurningRoutePedestrian, the bicycle moves faster and the ego has to brake earlier.
4. BlockedIntersection
While performing a maneuver, the ego vehicle encounters a stopped vehicle on the road and must perform an emergency brake or an avoidance maneuver.
5. SignalizedJunctionLeftTurn
The ego vehicle is performing an unprotected left turn at an intersection, yielding to oncoming traffic.
6. SignalizedJunctionLeftTurnEnterFlow
The ego vehicle is performing an unprotected left turn at an intersection, merging into opposite traffic.
7. NonSignalizedJunctionLeftTurn
Non-signalized version of SignalizedJunctionLeftTurn. The ego has to negotiate with the opposite vehicles without traffic lights.
8. NonSignalizedJunctionLeftTurnEnterFlow
Non-signalized version of SignalizedJunctionLeftTurnEnterFlow.
9. SignalizedJunctionRightTurn
The ego vehicle is turning right at an intersection and has to safely merge into the traffic flow coming from its left.
10. NonSignalizedJunctionRightTurn
Non-signalized version of SignalizedJunctionRightTurn. The ego has to negotiate with the traffic flow without traffic lights.
11. EnterActorFlows
A flow of cars runs a red light in front of the ego when it enters the junction, forcing it to react (interrupting the flow or merging into the flow). These vehicles are ’special’ ones such as police cars, ambulances, or firetrucks.
12. HighwayExit
The ego vehicle must cross a lane of moving traffic to exit the highway at an off-ramp.
13. MergerIntoSlowTraffic
The ego vehicle must merge into a slow traffic flow on the off-ramp when exiting the highway.
14. MergerIntoSlowTrafficV2
The ego vehicle must merge into a slow traffic flow coming from the on-ramp when driving on highway roads.
15. InterurbanActorFlow
The ego vehicle leaves the interurban road by turning left, crossing a fast traffic flow.
16. InterurbanAdvancedActorFlow
The ego vehicle incorporates into the interurban road by turning left, first crossing a fast traffic flow, and then merging into another one.
17. HighwayCutIn
The ego vehicle encounters a vehicle merging into its lane from a highway on-ramp. The ego vehicle must decelerate, brake, or change lanes to avoid a collision.
18. CrossingBicycleFlow
The ego vehicle needs to perform a turn at an intersection yielding to bicycles crossing from either the left.
19. OppositeVehicleRunningRedLight
The ego vehicle is going straight at an intersection but a crossing vehicle runs a red light, forcing the ego vehicle to avoid the collision.
20. OppositeVehicleTakingPriority
Non-signalized version of OppositeVehicleTakingPriority.
21. VinillaTurn
A basic scenario for the ego vehicle to learn the basic traffic rules, e.g. stop signs and traffic lights.
Appendix 0.B Input & Output Representation
$C=34$
Static($C_{s}=1$) Dynamic($C_{d}=4$)
road route ego lane yellow line white line vehicle walker emergency car obstacle green traffic light yellow&red traffic light stop sign
Table 5: Dimension of input and output representation. The $T$ temporal masks are from historical time-steps: $[-16,-11,-6,-1]$.
H W C T M
128 128 34 4 30
@@ H W C T M
Throttle Brake Steer Throttle Brake Steer Throttle Brake Steer
@@
Appendix 0.C Simulator Execution Mode
We boost CARLA running efficiency via asynchronous reloading and parallel execution which could avoid the waiting time caused by the long preparation time of starting a new route, as shown in Fig. 4. For each 100K steps of execution, it can reduce about 1 day time cost.